x+ 5x-2/9-7=3x+4/3
Answer by jim_thompson5910(35256)

I think I answered this before (or already). If not, then here it is again.


Multiply everything by the LCD 9 to clear out the fractions to get


9x + 9(5x-2)/9 - 7*9 = 9(3x+4)/3

9x + 1(5x-2) - 63 = 3(3x+4)


Now solve for x


9x + 1(5x-2) - 63 = 3(3x+4)


9x + 1(5x)+1(-2) - 63 = 3(3x)+3(4)


9x + 5x - 2 - 63= 9x + 12


14x - 65 = 9x + 12


14x - 65 - 9x = 12


14x - 9x = 12 + 65


5x = 77


x = 77/5


So the solution is x = 77/5

Feel free to ask any questions you have. Thanks.